OpenStudy (anonymous):

Find the equation of the line that is perpendicular to 2x + 5y = 8 and passes through the point (−1, 0). show steps please Find the equation of the line that is perpendicular to 2x + 5y = 8 and passes through the point (−1, 0). show steps please @Mathematics

OpenStudy (aroub):

5y=-2x+8 y=-2/5x+8/5 slope=5/2 y=mx+b , m=5/2 , (-1,0) 0=5/2(-1)+b 0=-5/2+b 5/2=b so the equation is, y=5/2x+5/2